# Counting collisions in random circuit sampling for benchmarking quantum computers

### v1

@@ VII Conclusions and outlook
From the resource analysis presented in Sec. VI, one can envisage that the approach presented in this work can be applied in two different contexts: (i) for benchmarking small-scale quantum computers and (ii) for the validation (or cross-validation) of large-scale quantum computations, avoiding intensive classical computing costs and the associated energy consumption. Due to the exponential scaling of the sampling cost, the second application is not currently feasible for circuits which are hard to classically simulate ($n>50$ qubits), but may become feasible in the future if the sampling rate and/or the state fidelity will increase by one or two orders of magnitude. The first application is instead immediately feasible and can be easily used to benchmark existing quantum processors.
@@
  • The collision volume (CV) test, if applied on a single device, can be easily decieved by a “dishonest” machine. For example, a trivial classical device which always outputs the same bitstring would artificially pass the CV test. The cross-collision volume (XCV) test defined in Sec. IV seems much more difficult to spoof, unless both Alice and Bob agree on a cheating strategy. Is it possible to give a rigorous proof of this fact? Can Alice verify, by running a cross-collision test, that Bob has a good quantum computer without trusting him?
  • In this work we assumed the possibility of preparing Haar-random quantum states. In practice, however, one can only run random circuits of polynomial depth, which are known to generate good but not perfect approximations of Haar-random states. What happens to the results derived in this work if we consider feasible random circuits instead of ideal Haar-random states?
@@
  • Can one extend our benchmarks beyond quantum computers that are based on the circuit model? For example, one could imagine to benchmark analogue quantum devices, quantum simulators, continuous-variable photonic processors, etc., by running a protocol similar to Algorithm 1 or 2 but where the random circuit $C$ is replaced by some suitable state preparation procedure that approximately generates a Haar-random state.
@@ IX Acknowledgments
We thank the organizers and the participants of the NISQAH-2023 workshop in which some aspects of this work have been informally shared and discussed. In particular, we thank fruitful discussions with Scott Aaronson, who suggested to estimate the sampling cost in the quantum supremacy regime. We thank Nathan Shammah, William Zeng and Vincent Russo for useful comments and suggestions on this manuscript. AM acknowledges support from Unitary Fund and from the PNRR MUR project PE0000023-NQSTI (Italy).

### v2

@@ VII Conclusions and outlook
From the resource analysis presented in Sec. VI, one can envisage that the approach presented in this work can be applied in two different contexts: (i) for benchmarking small-scale quantum computers and (ii) for the validation (or cross-validation) of large-scale quantum computations, avoiding intensive classical computing costs and the associated energy consumption. Due to the exponential scaling of the sampling cost, the second application is not currently feasible for circuits that are hard to classically simulate ($n>50$ qubits), but may become feasible in the future if the sampling rate and/or the state fidelity will increase by one or two orders of magnitude. The first application is instead immediately feasible and can be easily used to benchmark existing quantum processors.
@@
  1. 1.
    The collision volume (CV) test, if applied on a single device, can be easily deceived by a “dishonest” machine. For example, a trivial classical device that always outputs the same bitstring would artificially pass the CV test. The cross-collision volume (XCV) test defined in Sec. IV seems much more difficult to spoof, unless both Alice and Bob agree on a cheating strategy. Is it possible to give a rigorous proof of this fact? Can Alice rigorously certify, by running a cross-collision test, that Bob has a good quantum computer without trusting him?
  2. 2.
    How sensitive are collision benchmarks to different types of errors? For example, systematic coherent errors such as miscalibration of gates are undetectable by counting collisions on a single device, since a wrong pure random state gives the same collision statistics as the correct pure random state. On the other hand, coherent errors are detectable by the cross-collision test applied on two independent devices, since different random pure states yield a small cross-collision anomaly (see Fig. 4). This fact suggests that, by comparing the number of local and cross-collisions, one may even deduce the relative impact between coherent and incoherent errors.
  3. 3.
    In this work we assumed the possibility of preparing Haar-random quantum states. In practice, however, one can only run random circuits of polynomial depth, which are known to generate good but not perfect approximations of Haar-random states. What happens to the results derived in this work if we consider feasible random circuits instead of ideal Haar-random states? On a similar vein, can we safely replace the Haar-random distribution with a unitary $t$-design?
  4. 4.
    Is it possible to apply similar benchmarks to random states that are subject to constraints, e.g. symmetries, in order to have a more peaked output distribution and therefore a larger quantum collision anomaly? Taking this question to the extreme limit, one can consider the recently proposed class of randomized circuits having the output distribution highly peaked on a specific random bitstring aaronson2024verifiable . What is the physical meaning of collisions and cross-collisions in this case? For a small number of samples, most collisions will trivially come from the peak bitstring, but what happens if we increase the number of shots?
  5. 5.
@@
  6. 6.
    Is there any workaround to reduce the high sampling cost of collision benchmarks? This seems an unavoidable limitation due to the small collision probability. Symmetries and more structured circuits (see point 4) may help reduce the sampling cost but, at the same time, could make classical spoofing easier. Moreover, a brute-force way of reducing the wall-clock time of the experiment is exploiting multi-core quantum processors as discussed in Sec. VI.
  7. 7.
    Can one extend our benchmarks beyond quantum computers that are based on the circuit model? For example, one could imagine benchmarking analog quantum devices, quantum simulators, continuous-variable photonic processors, etc., by running a protocol similar to Algorithm 1 or 2 but where the random circuit $C$ is replaced by some suitable state preparation procedure that approximately generates a Haar-random state.
  8. 8.
    In this work we always assumed a register of $n$ qubits. Can one extend the same results to $n$ qudits? Since we only used properties of Haar-random states in a global Hilbert space, we envisage that all results could be straightforwardly extended to qudits, up to minor adjustments, e.g., $D:=2^{n}\rightarrow d^{n}$, and bitstrings $\rightarrow$ ditstrings.
@@ IX Acknowledgments
We thank the organizers and the participants of the NISQAH-2023 workshop in which some aspects of this work have been informally shared and discussed. In particular, we thank fruitful discussions with Scott Aaronson, who suggested to estimate the sampling cost in the quantum supremacy regime. We thank Nathan Shammah, William Zeng and Vincent Russo for useful comments and suggestions on this manuscript. We acknowledge support from Unitary Fund and from the PNRR MUR project PE0000023-NQSTI (Italy).